# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v2

@@ 6.2 Proof of Lemma 4
Suppose that the rows of $\mathbf{F}_{\Omega}$ are chosen uniformly at random without replacement from $\mathbf{F}$, where $\mathbf{F^{*}F}=\mathbf{F}\mathbf{F}^{*}=\frac{N}{L}$. For a realization of $\mathbf{\mathbf{\Phi}}_{S}$ with $\|\mathbf{\mathbf{\Phi}}_{S}^{*}\mathbf{\mathbf{\Phi}}_{S}-\mathbf{I}_{S}\|\leq\eta_{1}$,
$\|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}-\mathbf{\mathbf{\Phi}}_{S}^{*}\mathbf{\mathbf{\Phi}}_{S}\|\leq\eta,\,\text{with probability at least}\,1-\epsilon$
@@
where $\sigma=\|\sum\limits_{l=1}^{L}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}^{1/2}\leq\max\limits_{l}\{|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}|\}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}=\|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}\|_{1\rightarrow 2}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}$. The derivation of the bound can be found, e.g. in Proposition 8.20 of [foucart_rauhut_2015].
By Proposition 7.13 in [foucart_rauhut_2015] regarding the moment estimation,
@@
Apply Lemma 14.3 [foucart_rauhut_2015] on the non-square matrix $\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}$, for any $p\geq 1$,
@@
Then, we apply a tail bound from moment estimation, see Proposition 7.15 from [foucart_rauhut_2015]. For $u\geq 1$,
@@
This completes the proof for the Bernoulli case. For the case when the rows of $\textbf{F}_{\Omega}$ are chosen uniformly from $F$, use a similar argument as in the later part of the proof of Theorem 14.1 [foucart_rauhut_2015], where it shows the probability for the uniform model is bounded by twice the one for the Bernoulli model derived here.
@@ 6.4 Proof of Lemma 7
Let $\textbf{V}=\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\textbf{X}_{0})$, define
$\boldsymbol{\mathcal{Y}}:=\textbf{A}^{*}\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\text{sgn}(\textbf{X}_{0}).$
$\boldsymbol{\mathcal{Y}}$ is the exact dual certificate since $\mathcal{P}_{S}\boldsymbol{\mathcal{Y}}=\text{sgn}(\textbf{X}_{0})$. Next, we have
$\begin{split}\|\textbf{V}\|_{F}&=\|\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}sgn(\textbf{X}_{0})\|_{F}\leq\frac{\sqrt{1+\eta}}{1-\eta}\sqrt{n}.\end{split}$ (41)
@@ 6.5 Proof of Lemma 8
Define $\tilde{X}$$=\begin{bmatrix}\begin{array}[]{c | c | c | c}\mathbf{h_{1}}sgn(z_{1})&\mathbf{h_{2}}sgn(z_{1})&\cdots&\mathbf{h_{c}}sgn(z_{1})\\ \hline\cr\mathbf{h_{1}}sgn(z_{2})&\mathbf{h_{2}}sgn(z_{2})&\cdots&\mathbf{h_{c}}sgn(z_{2})\\ \hline\cr\vdots&\vdots&\vdots\\ \mathbf{h_{1}}sgn(z_{N})&\mathbf{h_{2}}sgn(z_{N})&\cdots&\mathbf{h_{c}}sgn(z_{N})\end{array}\end{bmatrix}$. Then $sgn(\textbf{X}_{0})=$ $\tilde{X}$$/\|\textbf{H}\|_{F}\in\mathbb{C}^{kN\times C}$ , where $\mathbf{H}=\begin{bmatrix}\mathbf{h_{1}}|\mathbf{h_{2}}|\cdots|\mathbf{h_{C}}\end{bmatrix}$, $\mathbf{H}\in\mathbb{C}^{k\times C}.$
We want to estimate the norm of $\max\limits_{l\in S^{c}}\|\operatorname{sgn}(\textbf{X}_{0})^{*}\textbf{A}^{\dagger}\textbf{A}_{S^{l}}\|_{F}$,
@@
By the Cauchy-Schwartz inequality, for any y, $\|$$\tilde{X}$$\textbf{y}\|\leq\|\textbf{H}\|_{2\rightarrow 2}\|\operatorname{sgn}(\textbf{z})\|_{2}\|\textbf{y}\|_{2}\leq\sqrt{n}\|\textbf{H}\|_{2\rightarrow 2}$.
We want to bound the two terms, $\|\textbf{H}\|_{F}$ and $\|\textbf{H}\|_{2\rightarrow 2}$. For $\|\textbf{H}\|_{F}$, since each column of H has norm $1$, we have $\|\textbf{H}\|_{F}=\sqrt{C}$.
To bound $\|\textbf{H}\|_{2\rightarrow 2}$, we apply the matrix Bernstein inequality to $\textbf{H}^{*}\textbf{H}=\Sigma_{l=1}^{C}\textbf{h}_{l}\textbf{h}_{l}^{*}$. Define $\textbf{Z}_{l}=\textbf{h}_{l}\textbf{h}_{l}^{*}-\frac{1}{k}\textbf{I}_{k}$, $\textbf{Z}_{l}$ are independent with $\mathbb{E}(\textbf{Z}_{l})=0$,
$R=\max_{l}\|\textbf{Z}_{l}\|=\|\textbf{h}_{l}\textbf{h}_{l}^{*}-\frac{1}{k}I\|\leq\max_{l}\max(\|\textbf{h}_{l}\|^{2},\frac{1}{k})=1,$ (42)
$\sigma^{2}=\|\Sigma_{l=1}^{C}\mathbb{E}(\textbf{Z}_{l}\textbf{Z}_{l}^{*})\|=\|\Sigma_{l=1}^{C}\mathbb{E}(\textbf{h}_{l}\underbrace{\textbf{h}_{l}^{*}\textbf{h}_{l}}_{=1}\textbf{h}_{l}^{*})-\frac{2}{k}\mathbb{E}(\textbf{h}_{l}\textbf{h}_{l}^{*})+\frac{I}{k^{2}}\|=\|\Sigma_{l=1}^{C}(\frac{I}{k}-\frac{I}{k^{2}})\|\leq C/k.$ (43)
Hence,
$\mathbb{P}(\|\textbf{H}\textbf{H}^{*}-\frac{C}{k}I\|\geq t)\leq 2k\exp(-\frac{t^{2}/2}{\sigma^{2}+Rt/3})=2k\exp(-\frac{t^{2}/2}{\frac{C}{k}+t/3}).$ (44)
Take $t=\frac{C}{2k}$, we have $\|\textbf{H}\|\leq\sqrt{3C/2k}$ with probability at least $1-2k\exp(-\frac{3C}{28k})$.
Combining the bounds, we have
$\frac{\sqrt{1+\eta}}{1-\eta}\sqrt{\frac{1}{L}}\mu^{\mathbf{\mathbf{\mathbf{\mathbf{\Psi}}}}}_{max}\sqrt{3n/2}<\theta<1$ (45)
@@
The case when the coefficients $\textbf{h}_{l},l\in[C]$ are chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$ follows the exact argument by noticing that $\textbf{h}_{l}=\frac{\textbf{s}_{l}}{\|\textbf{s}_{l}\|_{2}}$ where $\textbf{s}_{l}$ are independent Gaussian vectors drawn from $\frac{1}{\sqrt{2}}\mathcal{N}(0,1)+i\frac{1}{\sqrt{2}}\mathcal{N}(0,1)$. $\square$
Acknowledgement
@@ 7.1 Conditioning of random submatrices
Then, with probability at least ($1-\epsilon$), $\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|\leq\eta.$
@@ 7.2 Algorithm
Recall from (26), $\|\mathbf{X}\|_{1,2}=\sum\limits_{j\in[N]}\|\mathbf{X}_{T_{j}\times[C]}\|_{F}$. Let $f(\mathbf{X})=\frac{1}{2}\|\textbf{Y}-\textbf{A}\textbf{X}\|_{F}^{2}$ and $g(\mathbf{X})=\lambda\|\textbf{X}\|_{1,2}$, an basic update of ISTA [Bubeck_2009] for problem (29) at iteration $k$ using stepsize $L$ is,

### v3

@@ 6.2 Proof of Lemma 5
Suppose that the rows of $\mathbf{F}_{\Omega}$ are chosen uniformly at random without replacement from $\mathbf{F}$, where $\mathbf{F^{*}F}=\mathbf{F}\mathbf{F}^{*}=\frac{N}{L}$. For a realization of $\mathbf{\mathbf{\Phi}}_{S}$ with $\|\mathbf{\mathbf{\Phi}}_{S}^{*}\mathbf{\mathbf{\Phi}}_{S}-\mathbf{I}_{S}\|_{2\rightarrow 2}\leq\eta_{1}$,
$\|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}-\mathbf{\mathbf{\Phi}}_{S}^{*}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}\leq\eta,\,\text{with probability at least}\,1-\epsilon$
@@
where $\sigma=\|\sum\limits_{l=1}^{L}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}\textbf{a}_{l}^{*}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}^{1/2}\leq\max\limits_{l}\{|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{a}_{l}|\}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}=\|\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}\|_{1\rightarrow 2}\|\textbf{F}_{\Omega}\mathbf{\mathbf{\Phi}}_{S}\|_{2\rightarrow 2}$. The derivation of the bound can be found, e.g. in Proposition 8.20 of [21].
By Proposition 7.13 in [21] regarding the moment estimation,
@@
Apply Lemma 14.3 [21] on the non-square matrix $\mathbf{\mathbf{\Phi}}_{S}^{*}\textbf{F}_{\Omega}^{*}$, for any $p\geq 1$,
@@
Then, we apply a tail bound from moment estimation, see Proposition 7.15 from [21]. For $u\geq 1$,
@@
This completes the proof for the Bernoulli case. For the case when the rows of $\textbf{F}_{\Omega}$ are chosen uniformly from $F$, use a similar argument as in the later part of the proof of Theorem 14.1 [21], where it shows the probability for the uniform model is bounded by twice the one for the Bernoulli model derived here.
@@ 6.4 Proof of Lemma 7
Let $\textbf{V}=\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\operatorname{sgn}(\textbf{X}_{0})$, define
$\boldsymbol{\mathcal{Y}}:=\textbf{A}^{*}\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\operatorname{sgn}(\textbf{X}_{0}).$
$\boldsymbol{\mathcal{Y}}$ is the exact dual certificate since $\mathcal{P}_{S}\boldsymbol{\mathcal{Y}}=\operatorname{sgn}(\textbf{X}_{0})$. Next, we have
$\begin{split}\|\textbf{V}\|_{F}&=\|\textbf{A}_{S}(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\operatorname{sgn}(\textbf{X}_{0})\|_{F}\leq\frac{\sqrt{1+\eta}}{1-\eta}\sqrt{n}.\end{split}$ (41)
@@ 6.5 Proof of Lemma 8
Define $\tilde{X}$$\begin{bmatrix}\begin{array}[]{c | c | c | c}\mathbf{h_{1}}\operatorname{sgn}(z_{1})&\mathbf{h_{2}}\operatorname{sgn}(z_{1})&\cdots&\mathbf{h_{c}}\operatorname{sgn}(z_{1})\\ \hline\cr\mathbf{h_{1}}\operatorname{sgn}(z_{2})&\mathbf{h_{2}}\operatorname{sgn}(z_{2})&\cdots&\mathbf{h_{c}}\operatorname{sgn}(z_{2})\\ \hline\cr\vdots&\vdots&\vdots\\ \mathbf{h_{1}}\operatorname{sgn}(z_{N})&\mathbf{h_{2}}\operatorname{sgn}(z_{N})&\cdots&\mathbf{h_{c}}\operatorname{sgn}(z_{N})\end{array}\end{bmatrix}$. Then $\operatorname{sgn}(\textbf{X}_{0})=$ $\tilde{X}$$\|\textbf{H}\|_{F}\in\mathbb{C}^{kN\times C}$ , where $\mathbf{H}=\begin{bmatrix}\mathbf{h_{1}}|\mathbf{h_{2}}|\cdots|\mathbf{h_{C}}\end{bmatrix}$, $\mathbf{H}\in\mathbb{C}^{k\times C}.$
First, we estimate the norm of $\max\limits_{l\in S^{c}}\|\operatorname{sgn}(\textbf{X}_{0})^{*}\textbf{A}^{\dagger}\textbf{A}_{S^{l}}\|_{F}$:
@@
[…]
To bound $\|\textbf{H}\|_{2\rightarrow 2}$, we apply the non-commutative matrix Bernstein inequality (Theorem 4.4 in [1]) to $\textbf{H}\textbf{H}^{*}=\Sigma_{l=1}^{C}\textbf{h}_{l}\textbf{h}_{l}^{*}$. Consider the independent and centered random matrices $\textbf{Z}_{l}=\textbf{h}_{l}\textbf{h}_{l}^{*}-\frac{1}{k}\textbf{I}_{k}$ with $\mathbb{E}(\textbf{Z}_{l})=0$ and dimension $k\times k$,
$R=\max_{l}\|\textbf{Z}_{l}\|=\|\textbf{h}_{l}\textbf{h}_{l}^{*}-\frac{1}{k}\mathbf{I}\|\leq\max_{l}\max\big(\|\textbf{h}_{l}\|^{2},\frac{1}{k}\big)=1,$ (42)
$\sigma^{2}=\max\{\|\sum\limits_{l=1}^{C}\mathbb{E}(\textbf{Z}_{l}\textbf{Z}_{l}^{*})\|,\|\sum\limits_{l=1}^{C}\mathbb{E}(\textbf{Z}_{l}^{*}\textbf{Z}_{l})\|\}=\|\sum\limits_{l=1}^{C}\mathbb{E}(\textbf{h}_{l}\underbrace{\textbf{h}_{l}^{*}\textbf{h}_{l}}_{=1}\textbf{h}_{l}^{*})-\frac{2}{k}\mathbb{E}(\textbf{h}_{l}\textbf{h}_{l}^{*})+\frac{I}{k^{2}}\|=\|\sum\limits_{l=1}^{C}(\frac{I}{k}-\frac{I}{k^{2}})\|\leq C/k.$ (43)
[…]
$\mathbb{P}(\|\textbf{H}\textbf{H}^{*}-\frac{C}{k}\textbf{I}\|\geq t)\leq 2k\exp\Big(-\frac{t^{2}/2}{\sigma^{2}+Rt/3}\Big)=2k\exp\Big(-\frac{t^{2}/2}{\frac{C}{k}+t/3}\Big).$ (44)
Take $t=\frac{C}{2k}$, we have $\|\textbf{H}\|\leq\sqrt{3C/2k}$ with probability at least $1-2k\exp\big(-\frac{3C}{28k}\big)$.
Lastly, we estimate the $\alpha$ in the bound of $\max\limits_{j\in S^{c}}\|\textbf{A}^{\dagger}\textbf{A}_{S_{j}}\|_{F}\leq\alpha$. First, conditioned on $\|\mathbf{A}^{*}_{S}\mathbf{A}_{S}-\mathbf{I}_{S}\|\leq\eta<1,$
$\begin{split}\max\limits_{j\in S^{c}}\|\textbf{A}^{\dagger}\textbf{A}_{S_{j}}\|_{F}\leq\|(\textbf{A}_{S}^{*}\textbf{A}_{S})^{-1}\|_{2\rightarrow 2}\|\textbf{A}_{S}\|_{2\rightarrow 2}\max\limits_{j\in S^{c}}\|\textbf{A}_{S_{j}}\|_{F}\leq\frac{\sqrt{1+\eta}}{1-\eta}\max\limits_{j\in S^{c}}\|\textbf{A}_{S_{j}}\|_{F}.\end{split}$
[…]
Combining the above bounds, we have
$\max\limits_{j\in S^{c}}\|\textbf{A}^{\dagger}\textbf{A}_{S_{j}}\|_{F}\leq\frac{\sqrt{1+\eta}}{1-\eta}\sqrt{\frac{k}{L}}\mu^{\Psi}_{max}.$
Hence, when $L>6{\mu^{\Psi}_{max}}^{2}\frac{1+\eta}{(1-\eta)^{2}}n$, we have
$\max\limits_{l\in S^{c}}\|\operatorname{sgn}(\textbf{X}_{0})^{*}\textbf{A}^{\dagger}\textbf{A}_{S^{l}}\|_{F}\leq\frac{\sqrt{1+\eta}}{1-\eta}\sqrt{\frac{1}{L}}\mu^{\mathbf{\mathbf{\mathbf{\mathbf{\Psi}}}}}_{max}\sqrt{3n/2}<1/2$ (45)
@@
Equivalently, the result can be stated as $\max\limits_{j\in[|S^{c}|]}\|\mathcal{P}_{S^{c}_{j}}\boldsymbol{\mathcal{Y}}\|_{F}<1/2$ with probability at least $1-N^{-\alpha}$ if
$C\geq\tilde{C}_{\alpha}\log(N)k,$
where $\tilde{C}_{\alpha}$ grows linearly with respect to $\alpha$.
The case when the coefficients $\textbf{h}_{l},l\in[C]$ are chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$ follows an analogous argument by noticing that $\textbf{h}_{l}=\frac{\textbf{s}_{l}}{\|\textbf{s}_{l}\|_{2}}$ where $\textbf{s}_{l}$ are independent Gaussian vectors drawn from $\frac{1}{\sqrt{2}}\mathcal{N}(0,1)+i\frac{1}{\sqrt{2}}\mathcal{N}(0,1)$. $\square$
Acknowledgment
@@
Code Availability
The source code for this project will be made openly available at https://github.com/mathgirlstarttocode/AutoCalib_BiConvex_pMRI.git upon the acceptance of this work for publication.
@@ 7.1 Conditioning of random submatrices
Then, with probability at least ($1-\epsilon$), $\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|_{2\rightarrow 2}\leq\eta.$
@@ 7.2 Algorithm
Recall from (26), $\|\mathbf{X}\|_{1,2}=\sum\limits_{j\in[N]}\|\mathbf{X}_{T_{j}\times[C]}\|_{F}$. Let $f(\mathbf{X})=\frac{1}{2}\|\textbf{Y}-\textbf{A}\textbf{X}\|_{F}^{2}$ and $g(\mathbf{X})=\lambda\|\textbf{X}\|_{1,2}$, an basic update of ISTA [67] for problem (29) at iteration $k$ using stepsize $L$ is,